OpenStudy (jessicawade):

help me with mean and population standard distribution. WILL MEDAL AND FAN

OpenStudy (jessicawade):

Calculate the mean and the population standard deviation for the data set, rounded to thousandths. {2, 3, 5, 0, 3, 2, 1, 1, 5, 7, 8, 5, 3}

OpenStudy (anonymous):

mean=\[\mu=\frac{2+3+5+0+3+2+1+1+5+7+8+5+3}{number of elements=13}\]=\[\frac{45}{13}\] sd =\[\sqrt{var} \] = \[\sqrt{|each number - mean|^2} \] =\[\sqrt{\frac{{(|2-\mu|^2+|3-\mu|^2+|5-\mu|^2+|0-\mu|^2+|3-\mu|^2+|2-\mu|^2+|1-\mu|^2+|1-\mu|^2+|5-\mu|^2+|7-\mu|^2+|8-\mu|^2+|5-\mu|^2+|3-\mu|^2)}}{((13-1)}}\]=\[5\cdot\sqrt{\frac{3}{13}}\]

OpenStudy (jessicawade):

what lol

OpenStudy (jessicawade):

you lost me at var i havnt learned that one yet

OpenStudy (jessicawade):

the mean is 3.46

OpenStudy (jessicawade):

@Hero

OpenStudy (jessicawade):

@hartnn

OpenStudy (jessicawade):

@bakur

OpenStudy (anonymous):

313−−−√

OpenStudy (jessicawade):

whats 313?

OpenStudy (jessicawade):

@jmatthews4929

OpenStudy (anonymous):

yes, @jassicawade. \[45/13=3.46\], var = variance which is the how far each item in the population is to the mean \[\mu\], it is also defined as : average squared deviation from the mean.